OpenStudy (anonymous):

A company has two large computers. The slower computer can send all the company’s email in 40 minutes. The faster computer can complete the same job in 10 minutes. If both computers are working together, how long will it take them to do the job?

OpenStudy (aum):

\[ \frac 1x = \frac{1}{40} + \frac{1}{10} \]where x is the time it takes in minutes of both computers are working together. Solve for x.

OpenStudy (aum):

*if*